\begin{thebibliography}{75}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Bengio et~al.(2015)Bengio, Vinyals, Jaitly, and
  Shazeer}]{bengio2015scheduled}
Bengio, S.; Vinyals, O.; Jaitly, N.; and Shazeer, N. 2015.
\newblock Scheduled sampling for sequence prediction with recurrent neural
  networks.
\newblock \emph{Advances in neural information processing systems}, 28.

\bibitem[{Boult et~al.(2019)Boult, Cruz, Dhamija, Gunther, Henrydoss, and
  Scheirer}]{boult2019learning}
Boult, T.~E.; Cruz, S.; Dhamija, A.~R.; Gunther, M.; Henrydoss, J.; and
  Scheirer, W.~J. 2019.
\newblock Learning and the unknown: Surveying steps toward open world
  recognition.
\newblock In \emph{Proceedings of the AAAI conference on artificial
  intelligence}, volume~33, 9801--9807.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan,
  Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess,
  Clark, Berner, McCandlish, Radford, Sutskever, and
  Amodei}]{brown2020language}
Brown, T.; Mann, B.; Ryder, N.; Subbiah, M.; Kaplan, J.~D.; Dhariwal, P.;
  Neelakantan, A.; Shyam, P.; Sastry, G.; Askell, A.; Agarwal, S.;
  Herbert-Voss, A.; Krueger, G.; Henighan, T.; Child, R.; Ramesh, A.; Ziegler,
  D.; Wu, J.; Winter, C.; Hesse, C.; Chen, M.; Sigler, E.; Litwin, M.; Gray,
  S.; Chess, B.; Clark, J.; Berner, C.; McCandlish, S.; Radford, A.; Sutskever,
  I.; and Amodei, D. 2020.
\newblock Language Models are Few-Shot Learners.
\newblock In Larochelle, H.; Ranzato, M.; Hadsell, R.; Balcan, M.; and Lin, H.,
  eds., \emph{Advances in Neural Information Processing Systems}, volume~33,
  1877--1901. Curran Associates, Inc.

\bibitem[{Buzzega et~al.(2020)Buzzega, Boschini, Porrello, Abati, and
  CALDERARA}]{buzzega2020dark}
Buzzega, P.; Boschini, M.; Porrello, A.; Abati, D.; and CALDERARA, S. 2020.
\newblock Dark Experience for General Continual Learning: a Strong, Simple
  Baseline.
\newblock In Larochelle, H.; Ranzato, M.; Hadsell, R.; Balcan, M.; and Lin, H.,
  eds., \emph{Advances in Neural Information Processing Systems}, volume~33,
  15920--15930. Curran Associates, Inc.

\bibitem[{Chaudhry et~al.(2019{\natexlab{a}})Chaudhry, Ranzato, Rohrbach, and
  Elhoseiny}]{chaudhry2019efficient}
Chaudhry, A.; Ranzato, M.; Rohrbach, M.; and Elhoseiny, M. 2019{\natexlab{a}}.
\newblock Efficient lifelong learning with a-gem.
\newblock In \emph{Proceedings of ICLR}.

\bibitem[{Chaudhry et~al.(2019{\natexlab{b}})Chaudhry, Rohrbach, Elhoseiny,
  Ajanthan, Dokania, Torr, and Ranzato}]{chaudhry2019tiny}
Chaudhry, A.; Rohrbach, M.; Elhoseiny, M.; Ajanthan, T.; Dokania, P.~K.; Torr,
  P. H.~S.; and Ranzato, M. 2019{\natexlab{b}}.
\newblock On Tiny Episodic Memories in Continual Learning.
\newblock arXiv:1902.10486.

\bibitem[{Chen, Goodfellow, and Shlens(2016)}]{chen2016net2net}
Chen, T.; Goodfellow, I.; and Shlens, J. 2016.
\newblock {Net2Net}: Accelerating Learning via Knowledge Transfer.
\newblock In \emph{Proceedings of ICLR}.

\bibitem[{Dasigi et~al.(2019)Dasigi, Liu, Marasovi{\'c}, Smith, and
  Gardner}]{dasigi-etal-2019-quoref}
Dasigi, P.; Liu, N.~F.; Marasovi{\'c}, A.; Smith, N.~A.; and Gardner, M. 2019.
\newblock {Q}uoref: A Reading Comprehension Dataset with Questions Requiring
  Coreferential Reasoning.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, 5925--5932. Hong Kong, China:
  Association for Computational Linguistics.

\bibitem[{Dua et~al.(2019)Dua, Wang, Dasigi, Stanovsky, Singh, and
  Gardner}]{dua-etal-2019-drop}
Dua, D.; Wang, Y.; Dasigi, P.; Stanovsky, G.; Singh, S.; and Gardner, M. 2019.
\newblock {DROP}: A Reading Comprehension Benchmark Requiring Discrete
  Reasoning Over Paragraphs.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, 2368--2378. Minneapolis,
  Minnesota: Association for Computational Linguistics.

\bibitem[{Ebrahimi et~al.(2020)Ebrahimi, Meier, Calandra, Darrell, and
  Rohrbach}]{ebrahimi2020adversarial}
Ebrahimi, S.; Meier, F.; Calandra, R.; Darrell, T.; and Rohrbach, M. 2020.
\newblock Adversarial continual learning.
\newblock In \emph{European Conference on Computer Vision}, 386--402. Springer.

\bibitem[{Farajtabar et~al.(2020)Farajtabar, Azizan, Mott, and
  Li}]{pmlr-v108-farajtabar20a}
Farajtabar, M.; Azizan, N.; Mott, A.; and Li, A. 2020.
\newblock Orthogonal Gradient Descent for Continual Learning.
\newblock In Chiappa, S.; and Calandra, R., eds., \emph{Proceedings of the
  Twenty Third International Conference on Artificial Intelligence and
  Statistics}, volume 108 of \emph{Proceedings of Machine Learning Research},
  3762--3773. PMLR.

\bibitem[{Fernando et~al.(2017)Fernando, Banarse, Blundell, Zwols, Ha, Rusu,
  Pritzel, and Wierstra}]{fernando2017pathnet}
Fernando, C.; Banarse, D.; Blundell, C.; Zwols, Y.; Ha, D.; Rusu, A.~A.;
  Pritzel, A.; and Wierstra, D. 2017.
\newblock PathNet: Evolution Channels Gradient Descent in Super Neural
  Networks.
\newblock arXiv:1701.08734.

\bibitem[{French(1999)}]{french1999catastrophic}
French, R.~M. 1999.
\newblock Catastrophic forgetting in connectionist networks.
\newblock \emph{Trends in cognitive sciences}, 3(4): 128--135.

\bibitem[{Gao et~al.(2022)Gao, Li, Shan, Qu, Wang, and Zhang}]{gao2022forget}
Gao, J.; Li, J.; Shan, H.; Qu, Y.; Wang, J.~Z.; and Zhang, J. 2022.
\newblock Forget Less, Count Better: A Domain-Incremental Self-Distillation
  Learning Benchmark for Lifelong Crowd Counting.
\newblock arXiv:2205.03307.

\bibitem[{Geigle et~al.(2021)Geigle, Reimers, Rücklé, and
  Gurevych}]{geigle2021tweac}
Geigle, G.; Reimers, N.; Rücklé, A.; and Gurevych, I. 2021.
\newblock TWEAC: Transformer with Extendable QA Agent Classifiers.
\newblock arXiv:2104.07081.

\bibitem[{Ghifary et~al.(2015)Ghifary, Kleijn, Zhang, and
  Balduzzi}]{Ghifary_2015_ICCV}
Ghifary, M.; Kleijn, W.~B.; Zhang, M.; and Balduzzi, D. 2015.
\newblock Domain Generalization for Object Recognition With Multi-Task
  Autoencoders.
\newblock In \emph{Proceedings of the IEEE International Conference on Computer
  Vision (ICCV)}.

\bibitem[{Houlsby et~al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone,
  De~Laroussilhe, Gesmundo, Attariyan, and Gelly}]{pmlr-v97-houlsby19a}
Houlsby, N.; Giurgiu, A.; Jastrzebski, S.; Morrone, B.; De~Laroussilhe, Q.;
  Gesmundo, A.; Attariyan, M.; and Gelly, S. 2019.
\newblock Parameter-Efficient Transfer Learning for {NLP}.
\newblock In Chaudhuri, K.; and Salakhutdinov, R., eds., \emph{Proceedings of
  the 36th International Conference on Machine Learning}, volume~97 of
  \emph{Proceedings of Machine Learning Research}, 2790--2799. PMLR.

\bibitem[{Jiang et~al.(2020)Jiang, Xu, Araki, and Neubig}]{jiang2020soft}
Jiang, Z.; Xu, F.~F.; Araki, J.; and Neubig, G. 2020.
\newblock {How Can We Know What Language Models Know?}
\newblock \emph{Transactions of the Association for Computational Linguistics},
  8: 423--438.

\bibitem[{Khashabi et~al.(2020)Khashabi, Min, Khot, Sabharwal, Tafjord, Clark,
  and Hajishirzi}]{khashabi-etal-2020-unifiedqa}
Khashabi, D.; Min, S.; Khot, T.; Sabharwal, A.; Tafjord, O.; Clark, P.; and
  Hajishirzi, H. 2020.
\newblock {UNIFIEDQA}: Crossing Format Boundaries with a Single {QA} System.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2020}, 1896--1907. Online: Association for Computational Linguistics.

\bibitem[{Khosla et~al.(2012)Khosla, Zhou, Malisiewicz, Efros, and
  Torralba}]{khosla2012undoing}
Khosla, A.; Zhou, T.; Malisiewicz, T.; Efros, A.~A.; and Torralba, A. 2012.
\newblock Undoing the Damage of Dataset Bias.
\newblock In Fitzgibbon, A.; Lazebnik, S.; Perona, P.; Sato, Y.; and Schmid,
  C., eds., \emph{Computer Vision -- ECCV 2012}, 158--171. Berlin, Heidelberg:
  Springer Berlin Heidelberg.
\newblock ISBN 978-3-642-33718-5.

\bibitem[{Kirkpatrick et~al.(2017)Kirkpatrick, Pascanu, Rabinowitz, Veness,
  Desjardins, Rusu, Milan, Quan, Ramalho, Grabska-Barwinska
  et~al.}]{kirkpatrick2017overcoming}
Kirkpatrick, J.; Pascanu, R.; Rabinowitz, N.; Veness, J.; Desjardins, G.; Rusu,
  A.~A.; Milan, K.; Quan, J.; Ramalho, T.; Grabska-Barwinska, A.; et~al. 2017.
\newblock Overcoming catastrophic forgetting in neural networks.
\newblock \emph{Proceedings of NAS}, 3521--3526.

\bibitem[{Kocisky et~al.(2018)Kocisky, Schwarz, Blunsom, Dyer, Hermann, Melis,
  and Grefenstette}]{TACL1197}
Kocisky, T.; Schwarz, J.; Blunsom, P.; Dyer, C.; Hermann, K.~M.; Melis, G.; and
  Grefenstette, E. 2018.
\newblock The NarrativeQA Reading Comprehension Challenge.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  6(0): 317--328.

\bibitem[{Kwiatkowski et~al.(2019)Kwiatkowski, Palomaki, Redfield, Collins,
  Parikh, Alberti, Epstein, Polosukhin, Devlin, Lee, Toutanova, Jones, Kelcey,
  Chang, Dai, Uszkoreit, Le, and Petrov}]{TACL1455}
Kwiatkowski, T.; Palomaki, J.; Redfield, O.; Collins, M.; Parikh, A.; Alberti,
  C.; Epstein, D.; Polosukhin, I.; Devlin, J.; Lee, K.; Toutanova, K.; Jones,
  L.; Kelcey, M.; Chang, M.-W.; Dai, A.; Uszkoreit, J.; Le, Q.; and Petrov, S.
  2019.
\newblock Natural Questions: a Benchmark for Question Answering Research.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  7(0): 452--466.

\bibitem[{Lai et~al.(2017)Lai, Xie, Liu, Yang, and Hovy}]{lai-etal-2017-race}
Lai, G.; Xie, Q.; Liu, H.; Yang, Y.; and Hovy, E. 2017.
\newblock {RACE}: Large-scale {R}e{A}ding Comprehension Dataset From
  Examinations.
\newblock In \emph{Proceedings of the 2017 Conference on Empirical Methods in
  Natural Language Processing}, 785--794. Copenhagen, Denmark: Association for
  Computational Linguistics.

\bibitem[{Lester, Al-Rfou, and Constant(2021)}]{lester-etal-2021-power}
Lester, B.; Al-Rfou, R.; and Constant, N. 2021.
\newblock The Power of Scale for Parameter-Efficient Prompt Tuning.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, 3045--3059. Online and Punta Cana, Dominican
  Republic: Association for Computational Linguistics.

\bibitem[{Li et~al.(2017)Li, Yang, Song, and Hospedales}]{Li2017deeper}
Li, D.; Yang, Y.; Song, Y.-Z.; and Hospedales, T.~M. 2017.
\newblock Deeper, Broader and Artier Domain Generalization.
\newblock In \emph{Proceedings of the IEEE International Conference on Computer
  Vision (ICCV)}.

\bibitem[{Li et~al.(2019)Li, Zhang, Yang, Liu, Song, and
  Hospedales}]{li2019episodic}
Li, D.; Zhang, J.; Yang, Y.; Liu, C.; Song, Y.-Z.; and Hospedales, T.~M. 2019.
\newblock Episodic training for domain generalization.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, 1446--1455.

\bibitem[{Li et~al.(2018)Li, Pan, Wang, and Kot}]{li2018domain}
Li, H.; Pan, S.~J.; Wang, S.; and Kot, A.~C. 2018.
\newblock Domain generalization with adversarial feature learning.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and
  pattern recognition}, 5400--5409.

\bibitem[{Li and Hoiem(2017)}]{li2017learning}
Li, Z.; and Hoiem, D. 2017.
\newblock Learning without forgetting.
\newblock \emph{TPAMI}, 40(12): 2935--2947.

\bibitem[{Lin(2004)}]{lin-2004-rouge}
Lin, C.-Y. 2004.
\newblock {ROUGE}: A Package for Automatic Evaluation of Summaries.
\newblock In \emph{Text Summarization Branches Out}, 74--81. Barcelona, Spain:
  Association for Computational Linguistics.

\bibitem[{Liu et~al.(2021)Liu, Zheng, Du, Ding, Qian, Yang, and
  Tang}]{liu2021gpt}
Liu, X.; Zheng, Y.; Du, Z.; Ding, M.; Qian, Y.; Yang, Z.; and Tang, J. 2021.
\newblock GPT Understands, Too.
\newblock arXiv:2103.10385.

\bibitem[{Loshchilov and Hutter(2017)}]{loshchilov2017decoupled}
Loshchilov, I.; and Hutter, F. 2017.
\newblock Decoupled Weight Decay Regularization.
\newblock arXiv:1711.05101.

\bibitem[{Madotto et~al.(2021)Madotto, Lin, Zhou, Moon, Crook, Liu, Yu, Cho,
  Fung, and Wang}]{madotto-etal-2021-continual}
Madotto, A.; Lin, Z.; Zhou, Z.; Moon, S.; Crook, P.; Liu, B.; Yu, Z.; Cho, E.;
  Fung, P.; and Wang, Z. 2021.
\newblock Continual Learning in Task-Oriented Dialogue Systems.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, 7452--7467. Online and Punta Cana, Dominican
  Republic: Association for Computational Linguistics.

\bibitem[{Mallya and Lazebnik(2018)}]{mallya2018packnet}
Mallya, A.; and Lazebnik, S. 2018.
\newblock Packnet: Adding multiple tasks to a single network by iterative
  pruning.
\newblock In \emph{Proceedings of the IEEE conference on Computer Vision and
  Pattern Recognition}, 7765--7773.

\bibitem[{Maltoni and Lomonaco(2019)}]{maltoni2019continuous}
Maltoni, D.; and Lomonaco, V. 2019.
\newblock Continuous learning in single-incremental-task scenarios.
\newblock \emph{Neural Networks}, 116: 56--73.

\bibitem[{Mancini et~al.(2018)Mancini, Ricci, Caputo, and
  Rota~Bulo}]{mancini2018adding}
Mancini, M.; Ricci, E.; Caputo, B.; and Rota~Bulo, S. 2018.
\newblock Adding new tasks to a single network with weight transformations
  using binary masks.
\newblock In \emph{Proceedings of the European Conference on Computer Vision
  (ECCV) Workshops}, 0--0.

\bibitem[{Mansoury et~al.(2020)Mansoury, Abdollahpouri, Pechenizkiy, Mobasher,
  and Burke}]{coverage}
Mansoury, M.; Abdollahpouri, H.; Pechenizkiy, M.; Mobasher, B.; and Burke, R.
  2020.
\newblock FairMatch: A Graph-Based Approach for Improving Aggregate Diversity
  in Recommender Systems.
\newblock In \emph{Proceedings of the 28th ACM Conference on User Modeling,
  Adaptation and Personalization}, UMAP '20, 154–162. New York, NY, USA:
  Association for Computing Machinery.
\newblock ISBN 9781450368612.

\bibitem[{McCann et~al.(2019)McCann, Keskar, Xiong, and Socher}]{mccann2019the}
McCann, B.; Keskar, N.~S.; Xiong, C.; and Socher, R. 2019.
\newblock The Natural Language Decathlon: Multitask Learning as Question
  Answering.

\bibitem[{Mi et~al.(2020)Mi, Chen, Zhao, Huang, and
  Faltings}]{mi-etal-2020-continual}
Mi, F.; Chen, L.; Zhao, M.; Huang, M.; and Faltings, B. 2020.
\newblock Continual Learning for Natural Language Generation in Task-oriented
  Dialog Systems.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2020}, 3461--3474. Online: Association for Computational Linguistics.

\bibitem[{Mihaylov et~al.(2018)Mihaylov, Clark, Khot, and
  Sabharwal}]{mihaylov-etal-2018-suit}
Mihaylov, T.; Clark, P.; Khot, T.; and Sabharwal, A. 2018.
\newblock Can a Suit of Armor Conduct Electricity? A New Dataset for Open Book
  Question Answering.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, 2381--2391. Brussels, Belgium: Association for
  Computational Linguistics.

\bibitem[{Muandet, Balduzzi, and Schölkopf(2013)}]{pmlr-v28-muandet13}
Muandet, K.; Balduzzi, D.; and Schölkopf, B. 2013.
\newblock Domain Generalization via Invariant Feature Representation.
\newblock In Dasgupta, S.; and McAllester, D., eds., \emph{Proceedings of the
  30th International Conference on Machine Learning}, volume~28 of
  \emph{Proceedings of Machine Learning Research}, 10--18. Atlanta, Georgia,
  USA: PMLR.

\bibitem[{Mundt et~al.(2020)Mundt, Hong, Pliushch, and
  Ramesh}]{mundt2020wholistic}
Mundt, M.; Hong, Y.~W.; Pliushch, I.; and Ramesh, V. 2020.
\newblock A Wholistic View of Continual Learning with Deep Neural Networks:
  Forgotten Lessons and the Bridge to Active and Open World Learning.
\newblock arXiv:2009.01797.

\bibitem[{Pu et~al.(2021)Pu, Chen, Liu, Bakker, and Lew}]{pu2021lifelong}
Pu, N.; Chen, W.; Liu, Y.; Bakker, E.~M.; and Lew, M.~S. 2021.
\newblock Lifelong person re-identification via adaptive knowledge
  accumulation.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition}, 7901--7910.

\bibitem[{Qin and Joty(2022)}]{qin2022lfpt}
Qin, C.; and Joty, S. 2022.
\newblock {LFPT}5: A Unified Framework for Lifelong Few-shot Language Learning
  Based on Prompt Tuning of T5.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{raffel2020exploring}
Raffel, C.; Shazeer, N.; Roberts, A.; Lee, K.; Narang, S.; Matena, M.; Zhou,
  Y.; Li, W.; and Liu, P.~J. 2020.
\newblock Exploring the Limits of Transfer Learning with a Unified Text-to-Text
  Transformer.
\newblock \emph{Journal of Machine Learning Research}, 21: 1--67.

\bibitem[{Rajpurkar et~al.(2016)Rajpurkar, Zhang, Lopyrev, and
  Liang}]{rajpurkar-etal-2016-squad}
Rajpurkar, P.; Zhang, J.; Lopyrev, K.; and Liang, P. 2016.
\newblock {SQ}u{AD}: 100,000+ Questions for Machine Comprehension of Text.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, 2383--2392. Austin, Texas: Association for
  Computational Linguistics.

\bibitem[{Rebuffi et~al.(2017)Rebuffi, Kolesnikov, Sperl, and
  Lampert}]{rebuffi2017icarl}
Rebuffi, S.-A.; Kolesnikov, A.; Sperl, G.; and Lampert, C.~H. 2017.
\newblock {iCaRL}: {I}ncremental classifier and representation learning.
\newblock In \emph{Proceedings of CVPR}, 2001--2010.

\bibitem[{Richardson, Burges, and Renshaw(2013)}]{richardson-etal-2013-mctest}
Richardson, M.; Burges, C.~J.; and Renshaw, E. 2013.
\newblock {MCT}est: A Challenge Dataset for the Open-Domain Machine
  Comprehension of Text.
\newblock In \emph{Proceedings of the 2013 Conference on Empirical Methods in
  Natural Language Processing}, 193--203. Seattle, Washington, USA: Association
  for Computational Linguistics.

\bibitem[{Ritter, Botev, and Barber(2018)}]{ritter2018online}
Ritter, H.; Botev, A.; and Barber, D. 2018.
\newblock Online structured laplace approximations for overcoming catastrophic
  forgetting.
\newblock In \emph{Proceedings of NIPS}, 3738--3748.

\bibitem[{Rogers, Gardner, and Augenstein(2021)}]{2107.12708}
Rogers, A.; Gardner, M.; and Augenstein, I. 2021.
\newblock QA Dataset Explosion: A Taxonomy of NLP Resources for Question
  Answering and Reading Comprehension.
\newblock arXiv:2107.12708.

\bibitem[{Rusu et~al.(2016)Rusu, Rabinowitz, Desjardins, Soyer, Kirkpatrick,
  Kavukcuoglu, Pascanu, and Hadsell}]{rusu2016progressive}
Rusu, A.~A.; Rabinowitz, N.~C.; Desjardins, G.; Soyer, H.; Kirkpatrick, J.;
  Kavukcuoglu, K.; Pascanu, R.; and Hadsell, R. 2016.
\newblock Progressive Neural Networks.
\newblock arXiv:1606.04671.

\bibitem[{Sap et~al.(2019)Sap, Rashkin, Chen, Le~Bras, and
  Choi}]{sap-etal-2019-social}
Sap, M.; Rashkin, H.; Chen, D.; Le~Bras, R.; and Choi, Y. 2019.
\newblock Social {IQ}a: Commonsense Reasoning about Social Interactions.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, 4463--4473. Hong Kong, China:
  Association for Computational Linguistics.

\bibitem[{Scellato et~al.(2010)Scellato, Mascolo, Musolesi, and
  Latora}]{scellato2010distance}
Scellato, S.; Mascolo, C.; Musolesi, M.; and Latora, V. 2010.
\newblock Distance matters: geo-social metrics for online social networks.
\newblock In \emph{3rd Workshop on Online Social Networks (WOSN 2010)}.

\bibitem[{Shin et~al.(2017)Shin, Lee, Kim, and Kim}]{shin2017continual}
Shin, H.; Lee, J.~K.; Kim, J.; and Kim, J. 2017.
\newblock Continual learning with deep generative replay.
\newblock In \emph{Proceedings of NIPS}, 2990--2999.

\bibitem[{Soares and Parreiras(2020)}]{soares2020literature}
Soares, M. A.~C.; and Parreiras, F.~S. 2020.
\newblock A literature review on question answering techniques, paradigms and
  systems.
\newblock \emph{Journal of King Saud University-Computer and Information
  Sciences}, 32(6): 635--646.

\bibitem[{Su et~al.(2020)Su, Guo, Zhang, Fan, Lan, and Cheng}]{su2020continual}
Su, L.; Guo, J.; Zhang, R.; Fan, Y.; Lan, Y.; and Cheng, X. 2020.
\newblock Continual domain adaptation for machine reading comprehension.
\newblock In \emph{Proceedings of the 29th ACM International Conference on
  Information \& Knowledge Management}, 1395--1404.

\bibitem[{Sun, Ho, and Lee(2019)}]{sun2019lamol}
Sun, F.-K.; Ho, C.-H.; and Lee, H.-Y. 2019.
\newblock LAMOL: LAnguage MOdeling for Lifelong Language Learning.
\newblock arXiv:1909.03329.

\bibitem[{Sun et~al.(2019)Sun, Yu, Chen, Yu, Choi, and Cardie}]{sun2019dream}
Sun, K.; Yu, D.; Chen, J.; Yu, D.; Choi, Y.; and Cardie, C. 2019.
\newblock Dream: A challenge data set and models for dialogue-based reading
  comprehension.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  7: 217--231.

\bibitem[{Thrun and Mitchell(1995)}]{thrun1995lifelong}
Thrun, S.; and Mitchell, T.~M. 1995.
\newblock Lifelong robot learning.
\newblock \emph{Robotics and autonomous systems}, 15(1-2): 25--46.

\bibitem[{Trischler et~al.(2017)Trischler, Wang, Yuan, Harris, Sordoni,
  Bachman, and Suleman}]{trischler-etal-2017-newsqa}
Trischler, A.; Wang, T.; Yuan, X.; Harris, J.; Sordoni, A.; Bachman, P.; and
  Suleman, K. 2017.
\newblock {N}ews{QA}: A Machine Comprehension Dataset.
\newblock In \emph{Proceedings of the 2nd Workshop on Representation Learning
  for {NLP}}, 191--200. Vancouver, Canada: Association for Computational
  Linguistics.

\bibitem[{Träuble et~al.(2022)Träuble, Goyal, Rahaman, Mozer, Kawaguchi,
  Bengio, and Schölkopf}]{2207.11240}
Träuble, F.; Goyal, A.; Rahaman, N.; Mozer, M.; Kawaguchi, K.; Bengio, Y.; and
  Schölkopf, B. 2022.
\newblock Discrete Key-Value Bottleneck.
\newblock arXiv:2207.11240.

\bibitem[{van~de Ven and Tolias(2019)}]{van2019three}
van~de Ven, G.~M.; and Tolias, A.~S. 2019.
\newblock Three scenarios for continual learning.
\newblock arXiv:1904.07734.

\bibitem[{Vu et~al.(2022)Vu, Lester, Constant, Al-Rfou{'}, and
  Cer}]{vu-etal-2022-spot}
Vu, T.; Lester, B.; Constant, N.; Al-Rfou{'}, R.; and Cer, D. 2022.
\newblock {SP}o{T}: Better Frozen Model Adaptation through Soft Prompt
  Transfer.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, 5039--5059. Dublin,
  Ireland: Association for Computational Linguistics.

\bibitem[{Wang et~al.(2022{\natexlab{a}})Wang, Zhang, Ebrahimi, Sun, Zhang,
  Lee, Ren, Su, Perot, Dy, and Pfister}]{wang2022dualprompt}
Wang, Z.; Zhang, Z.; Ebrahimi, S.; Sun, R.; Zhang, H.; Lee, C.-Y.; Ren, X.; Su,
  G.; Perot, V.; Dy, J.; and Pfister, T. 2022{\natexlab{a}}.
\newblock DualPrompt: Complementary Prompting for Rehearsal-free Continual
  Learning.
\newblock arXiv:2204.04799.

\bibitem[{Wang et~al.(2022{\natexlab{b}})Wang, Zhang, Lee, Zhang, Sun, Ren, Su,
  Perot, Dy, and Pfister}]{Wang_2022_CVPR}
Wang, Z.; Zhang, Z.; Lee, C.-Y.; Zhang, H.; Sun, R.; Ren, X.; Su, G.; Perot,
  V.; Dy, J.; and Pfister, T. 2022{\natexlab{b}}.
\newblock Learning To Prompt for Continual Learning.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition (CVPR)}, 139--149.

\bibitem[{Wiwatcharakoses and Berrar(2020)}]{chayut2020}
Wiwatcharakoses, C.; and Berrar, D. 2020.
\newblock SOINN+, a Self-Organizing Incremental Neural Network for Unsupervised
  Learning from Noisy Data Streams.
\newblock \emph{Expert Systems with Applications}, 143: 113069.

\bibitem[{Wortsman et~al.(2020)Wortsman, Ramanujan, Liu, Kembhavi, Rastegari,
  Yosinski, and Farhadi}]{wortsman2020sup}
Wortsman, M.; Ramanujan, V.; Liu, R.; Kembhavi, A.; Rastegari, M.; Yosinski,
  J.; and Farhadi, A. 2020.
\newblock Supermasks in Superposition.
\newblock In Larochelle, H.; Ranzato, M.; Hadsell, R.; Balcan, M.; and Lin, H.,
  eds., \emph{Advances in Neural Information Processing Systems}, volume~33,
  15173--15184. Curran Associates, Inc.

\bibitem[{Xu et~al.(2014)Xu, Li, Niu, and Xu}]{xu2014exploiting}
Xu, Z.; Li, W.; Niu, L.; and Xu, D. 2014.
\newblock Exploiting Low-Rank Structure from Latent Domains for Domain
  Generalization.
\newblock In Fleet, D.; Pajdla, T.; Schiele, B.; and Tuytelaars, T., eds.,
  \emph{Computer Vision -- ECCV 2014}, 628--643. Cham: Springer International
  Publishing.
\newblock ISBN 978-3-319-10578-9.

\bibitem[{Yang et~al.(2019)Yang, Xie, Lin, Li, Tan, Xiong, Li, and
  Lin}]{yang2019end}
Yang, W.; Xie, Y.; Lin, A.; Li, X.; Tan, L.; Xiong, K.; Li, M.; and Lin, J.
  2019.
\newblock End-to-End Open-Domain Question Answering with {BERT}serini.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics (Demonstrations)},
  72--77. Minneapolis, Minnesota: Association for Computational Linguistics.

\bibitem[{Ye et~al.(2021)Ye, Liu, Meng, and Li}]{biexp-ye}
Ye, H.; Liu, H.; Meng, F.; and Li, X. 2021.
\newblock Bi-Directional Exponential Angular Triplet Loss for RGB-Infrared
  Person Re-Identification.
\newblock \emph{IEEE Transactions on Image Processing}, 30: 1583--1595.

\bibitem[{Zenke, Poole, and Ganguli(2017)}]{zenke2017continual}
Zenke, F.; Poole, B.; and Ganguli, S. 2017.
\newblock Continual learning through synaptic intelligence.
\newblock In \emph{Proceedings of ICML}, 3987--3995.

\bibitem[{Zhang, Xu, and Lin(2021)}]{zhang2021deep}
Zhang, H.; Xu, H.; and Lin, T.-E. 2021.
\newblock Deep Open Intent Classification with Adaptive Decision Boundary.
\newblock In \emph{AAAI}, 14374--14382.

\bibitem[{Zhang et~al.(2020)Zhang, Zhang, Ghosh, Li, Tasci, Heck, Zhang, and
  Kuo}]{zhang2020class}
Zhang, J.; Zhang, J.; Ghosh, S.; Li, D.; Tasci, S.; Heck, L.; Zhang, H.; and
  Kuo, C.-C.~J. 2020.
\newblock Class-incremental learning via deep model consolidation.
\newblock In \emph{Proceedings of the IEEE/CVF Winter Conference on
  Applications of Computer Vision}, 1131--1140.

\bibitem[{Zhao et~al.(2020)Zhao, Gong, Liu, Fu, and Tao}]{zhao2020entropy}
Zhao, S.; Gong, M.; Liu, T.; Fu, H.; and Tao, D. 2020.
\newblock Domain Generalization via Entropy Regularization.
\newblock In Larochelle, H.; Ranzato, M.; Hadsell, R.; Balcan, M.; and Lin, H.,
  eds., \emph{Advances in Neural Information Processing Systems}, volume~33,
  16096--16107. Curran Associates, Inc.

\bibitem[{Zhong et~al.(2022)Zhong, Gao, Ding, Qin, Liu, Zhou, Wang, Yin, and
  Duan}]{zhong2022proqa}
Zhong, W.; Gao, Y.; Ding, N.; Qin, Y.; Liu, Z.; Zhou, M.; Wang, J.; Yin, J.;
  and Duan, N. 2022.
\newblock ProQA: Structural Prompt-based Pre-training for Unified Question
  Answering.
\newblock arXiv:2205.04040.

\end{thebibliography}

\documentclass[letterpaper]{article} 
\usepackage[submission]{aaai23}  
\usepackage{times}  
\usepackage{helvet}  
\usepackage{courier}  
\usepackage[hyphens]{url}  
\usepackage{graphicx} 
\usepackage{natbib}  
\usepackage{caption} 
\usepackage{algorithm}
\usepackage{algorithmic}
\usepackage{subfigure}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{booktabs}
\usepackage{multirow}
\usepackage{multicol}
\usepackage{makecell}
\usepackage{bm}

\newcolumntype{L}[1]{>{\raggedright\let\newline\\\arraybackslash\hspace{0pt}}m{#1}}
\newcolumntype{C}[1]{>{\centering\let\newline\\\arraybackslash\hspace{0pt}}m{#1}}
\newcolumntype{R}[1]{>{\raggedleft\let\newline\\\arraybackslash\hspace{0pt}}m{#1}}
\usepackage{newfloat}
\usepackage{listings}
\DeclareCaptionStyle{ruled}{labelfont=normalfont,labelsep=colon,strut=off} 
\floatstyle{ruled}
\newfloat{listing}{tb}{lst}{}
\floatname{listing}{Listing}
\title{Continuous Risk Prediction}
\author {
    Yi Dai
}

\begin{document}

\maketitle

\begin{abstract}
  Lifelong learning (LL) capabilities are essential for QA models to excel in real-world applications, and architecture-based LL approaches have proven to be a promising direction for achieving this goal. However, adapting existing methods to QA tasks is far from straightforward. Many prior approaches either rely on access to task identities during testing or fail to adequately model samples from unseen tasks, which limits their practical applicability. To overcome these limitations, we introduce Diana , a novel \underline{d}ynam\underline{i}c \underline{a}rchitecture-based lifelo\underline{n}g Q\underline{A} framework designed to learn a sequence of QA tasks using a prompt-enhanced language model.
  Diana leverages four hierarchically structured types of prompts to capture QA knowledge at multiple levels of granularity. Task-level prompts are specifically designed to encode task-specific knowledge, ensuring strong lifelong learning performance. Meanwhile, instance-level prompts are utilized to capture shared knowledge across diverse input samples, enhancing the model's generalization capabilities. Additionally, Diana incorporates dedicated prompts to explicitly handle unseen tasks and introduces a set of prompt key vectors that facilitate efficient knowledge transfer and sharing between tasks.
  Through extensive experimentation, we demonstrate that Diana achieves state-of-the-art performance among lifelong QA models, with particularly notable improvements in its ability to handle previously unseen tasks. This makes Diana a significant advancement in the field of lifelong learning for question-answering systems.
  \end{abstract}

\section{Introduction}
Question Answering (QA) plays a pivotal role in advancing AI applications \cite{yang2019end}. Modern QA models achieve remarkable performance when trained on fixed tasks \cite{soares2020literature}. However, these models often fall short in real-world scenarios where systems are expected to adapt to new QA tasks over time. Retraining a model on new tasks typically results in \emph{catastrophic forgetting}—a phenomenon where the model loses previously acquired knowledge while learning new information \cite{french1999catastrophic}. To address this issue without resorting to costly retraining, Lifelong Learning (LL) offers a solution by enabling models to continuously acquire new knowledge while preserving what they have already learned \cite{thrun1995lifelong}. Consequently, developing QA models with robust lifelong learning capabilities is critical for practical deployment.

One effective strategy for implementing LL is through architecture-based approaches \cite{chen2016net2net,rusu2016progressive,fernando2017pathnet,chayut2020,qin2022lfpt}, which maintain task-specific components to safeguard previously learned knowledge \cite{mancini2018adding}. Recently, researchers have explored converting QA tasks into a unified language modeling (LM) framework \cite{khashabi-etal-2020-unifiedqa}, achieving impressive results by leveraging prompts \cite{brown2020language,jiang2020soft} and adapter modules \cite{pmlr-v97-houlsby19a}. Some efforts have extended these techniques to create lifelong QA models by sharing a common pre-trained language model (PLM) across different tasks \cite{su2020continual}. These models demonstrate the ability to handle a sequence of QA tasks with varying formats \cite{zhong2022proqa}.

Despite these advancements, integrating lifelong learning into QA models presents two significant challenges:

\textbf{First}, many architecture-based LL methods require access to the task identity of test samples \cite{wang2022dualprompt,Wang_2022_CVPR}. This requirement severely limits their applicability in real-world settings, as task identities are often unavailable for input queries \cite{geigle2021tweac}. \textbf{Second}, most existing LL models are evaluated exclusively on seen tasks and fail to account for samples from unseen tasks \cite{boult2019learning,mundt2020wholistic}. This limitation is problematic in practical QA systems, where users frequently pose questions that do not align with previously encountered tasks.

To address these challenges, two main approaches have emerged:

\textbf{(1)} Task-level organization: Some models assign dedicated components to each task and use a classifier to predict the task identity for each test sample \cite{wortsman2020sup,madotto-etal-2021-continual}. While this approach achieves high LL performance when task identities are correctly predicted, it struggles with generalization in QA applications. Specifically, it becomes impractical to determine task identities for questions unrelated to previously seen tasks. Moreover, knowledge remains siloed within task-specific components, hindering the model's ability to generalize to unseen tasks \cite{2107.12708}.

\textbf{(2)} Instance-level organization: Other models adopt a dynamic architecture for each input sample, maintaining a pool of fine-grained components that are dynamically combined during inference based on the input instance \cite{chayut2020}. This approach eliminates the need to explicitly determine task identities for test samples \cite{Wang_2022_CVPR} and enhances the model's ability to handle unseen tasks by distributing knowledge across components \cite{2207.11240}. However, these models often yield suboptimal LL performance due to the absence of dedicated components for capturing task-specific knowledge \cite{wang2022dualprompt}.

In this work, we propose \textbf{Diana}, a novel \underline{d}ynam\underline{i}c \underline{a}rchitecture-based lifelo\underline{n}g Q\underline{A} model that combines the strengths of both approaches. Following prior work, we convert QA tasks into a unified LM format and leverage a prompt-enhanced PLM to learn these tasks. Diana maintains four types of hierarchically organized prompts—General Prompt, Format Prompt, Task Prompt, and Meta Prompt—to capture QA knowledge at varying granularities. Specifically, the \emph{general prompt} is shared across all tasks, while \emph{format prompts} are shared among tasks with the same QA format. A \emph{task prompt} is assigned to each incoming task, and a pool of \emph{meta prompts} is dynamically combined during inference. This dual-level organization allows Diana to achieve high LL performance while generalizing effectively to unseen tasks.

Furthermore, we introduce separate prompts for unseen tasks and learn a key vector for each task and meta prompt to facilitate knowledge sharing between tasks. This design enables explicit modeling of samples from unseen tasks, enhancing the model's adaptability.

We conduct extensive experiments on 11 benchmark QA tasks spanning three formats and evaluate the model's generalization capabilities on three unseen tasks. The results demonstrate that Diana surpasses state-of-the-art (SOTA) baselines across all benchmarks, particularly excelling in handling unseen tasks. Our key contributions are:

\begin{itemize}
  \item \textbf{Novel Framework}: We introduce Diana, a groundbreaking architecture-based lifelong QA model that employs four types of hierarchically organized prompts to capture knowledge at multiple levels. By maintaining both task-level and instance-level components, Diana achieves superior generalization to unseen tasks while maintaining high LL performance.
  
  \item \textbf{Explicit Handling of Unseen Tasks}: To the best of our knowledge, Diana is the first lifelong QA model to explicitly model unseen tasks. We allocate dedicated prompts for unseen tasks and introduce prompt keys to enhance cross-task knowledge sharing.

  \item \textbf{Empirical Validation}: Extensive experiments confirm that Diana outperforms SOTA baselines, establishing its effectiveness in building lifelong QA models.
\end{itemize}

This work represents a significant step forward in the development of practical, adaptable QA systems capable of thriving in dynamic, real-world environments.

\section{Related Work}
\subsection{Lifelong Learning}
Lifelong Learning aims to incrementally acquire new knowledge without forgetting previously learned information. Existing methods can be categorized into three main groups:

\textbf{Rehearsal-based methods} retain past knowledge by replaying samples from previous tasks \cite{rebuffi2017icarl,shin2017continual,sun2019lamol}.  
\textbf{Regularization-based methods} consolidate knowledge by introducing constraints on parameter updates \cite{kirkpatrick2017overcoming,zenke2017continual,li2017learning}.  
\textbf{Architecture-based methods} dynamically expand the model's capacity by adding task-specific components \cite{chen2016net2net,rusu2016progressive,fernando2017pathnet}.  

Despite their effectiveness, most existing LL methods assume access to task identities during testing, which limits their applicability in real-world QA systems.

\subsection{Unified Question Answering}
Recent studies have explored unifying QA tasks into a single framework \cite{khashabi-etal-2020-unifiedqa,mccann2019the}. For instance, ProQA \cite{zhong2022proqa} utilizes structured prompts to handle multiple QA formats. However, ProQA relies on task identities during testing and only considers a limited number of tasks. In contrast, Diana is designed to handle a broader range of tasks without requiring task identity annotations.

\section{Methodology}
\subsection{Task Formulation}
We aim to sequentially learn $N$ QA tasks $T_1, \cdots, T_N$, presented in $L$ different formats $F_1, \cdots, F_L$. Each task $T_i$ consists of training samples $(C, Q, A)$, where $C$ is the context, $Q$ is the question, and $A$ is the answer. Our model predicts $A$ based on $C$ and $Q$ using a prompt-enhanced encoder-decoder framework.

To evaluate generalization, we introduce $N'$ unseen tasks $T_{N+1}, \cdots, T_{N+N'}$, which are only used for testing. Importantly, task identities are not provided during testing.

\subsection{Hierarchical Prompt Structure}
Diana employs a hierarchical prompt structure to construct the input representation $P(C, Q)$:
\begin{equation}
P(C, Q) = [P_g; P_f(F_j); P_t(T_i); P_m(C, Q)],
\end{equation}
where:
- $P_g$: A general prompt shared across all tasks.
- $P_f(F_j)$: A format-specific prompt for tasks in format $F_j$.
- $P_t(T_i)$: A task-specific prompt for task $T_i$.
- $P_m(C, Q)$: A meta-prompt dynamically combined based on the input $(C, Q)$.

This structure enables Diana to capture both coarse-grained and fine-grained knowledge, enhancing its ability to generalize to unseen tasks.

\subsection{Key Vector Space Learning}
To infer task identities during testing, Diana associates each prompt with a key vector. Specifically:
- Task prompt keys $\bm{k}_t(T_i)$ are used to identify the most relevant task for a given input.
- Meta prompt keys $\bm{k}_m^j$ are used to dynamically combine meta-prompts based on similarity to the query vector $\bm{q} = h(C, Q)$.

We optimize key vectors using triplet loss to ensure that similar samples are grouped together while dissimilar ones are separated:
\begin{equation}
\mathcal{L}_t = \exp(||h(C, Q), \bm{k}_t(T_i)|| + \max(1 - ||h(C_n, Q_n), \bm{k}_t(T_i)||, 0)),
\end{equation}
where $(C_n, Q_n)$ is a negative sample selected from the memory buffer.

\subsection{Model Training}
Diana employs a two-stage training process:
1. **Key Vector Optimization**: Learn a representation space for prompt keys to facilitate task inference.
2. **Prompt and Model Optimization**: Fine-tune the constructed prompt $P(C, Q)$ and the backbone language model.

Additionally, we introduce a scheduled sampling mechanism to balance the use of ground truth and inferred task identities during training:
\begin{equation}
\epsilon_k = \max(0, \alpha - k\beta),
\end{equation}
where $\alpha$ and $\beta$ control the probability of using ground truth task identities.

\section{Experiments}
\subsection{Datasets and Metrics}
We evaluate Diana on 11 benchmark QA datasets spanning three formats: Extractive, Abstractive, and Multiple-Choice. The datasets include SQuAD \cite{rajpurkar-etal-2016-squad}, NewsQA \cite{trischler-etal-2017-newsqa}, and RACE \cite{lai-etal-2017-race}, among others. We reserve three datasets (Quoref, Drop, and Dream) as unseen tasks.

Performance is measured using:
- **Exact Match (EM)** for Extractive QA.
- **F1 Score** for Abstractive QA.
- **Accuracy** for Multiple-Choice QA.

\subsection{Results}
Table \ref{tab:seen} summarizes the results on seen tasks. Diana outperforms all baselines, achieving a relative improvement of 6.15\% on $A_N$ and a 27.26\% reduction in $F_N$ compared to the best baseline AFPER.

\begin{table}[t]
\centering
\small
\scalebox{0.9}{
\begin{tabular}{l|c|c|c}
\toprule
Methods & $A_N$ & $F_N$ & $A_{N'}$ \\
\midrule
AFPER & 52.69 & 9.28 & 36.79 \\
Diana w/o $\mathcal{M}$ & 50.30 & 12.68 & 39.22 \\
Diana & \textbf{55.93} & \textbf{6.75} & \textbf{42.12} \\
\bottomrule
\end{tabular}}
\caption{Performance comparison on seen tasks.}
\label{tab:seen}
\end{table}

\subsection{Analysis of Unseen Tasks}
On unseen tasks, Diana achieves a relative improvement of 9.49\% on $A_{N'}$ compared to the best baseline DER++. This demonstrates the effectiveness of our approach in handling novel tasks.

\section{Conclusion}
We introduced Diana, a novel lifelong QA model that leverages hierarchical prompts and dynamic task inference mechanisms. Experimental results show that Diana significantly outperforms state-of-the-art baselines, particularly in handling unseen tasks. Future work will explore extending Diana to other domains and modalities.
\bibliography{aaai23}
\end{document}